5m3 of a liquid whose density is 1.8 g / is going to be pumped from the floor level to a height of 12 m., If it is pumped in 45 min, what would be the pumping flow?
Select one:
to. 1211.11 l / m
b. 150.13 l / min
c. 160.13 l / min
d. 111.11 l / min
Answer #1

Ans.Total volume of liquid(V)=5m3 =5×1000= 5000 litres.(1 m​​​​​​3​​​​​ = 1000 litres)

Time taken in pumping the liquid(t) = 45 minutes

Volume flow rate or discharge (Q)= volume/time

Q= V/t = 5000/45= 111.11 l/min. Ans.( d)

Note. All given options are in l/min. So it is considered that volume flow rate is asked.